OpenStudy (anonymous):

evaluate the integral (36dx/36+e^x)

OpenStudy (amistre64):

k * integrate (k+e^x)^(-1) have you tried by parts?

OpenStudy (amistre64):

ln(k+e^x) e^x+k-k k --------- = 1 - ----- k+e^x k+e^x so if we add 1 and subtract 1, we should be able to work this

OpenStudy (anonymous):

\[\int\limits \frac{ 36 dx }{ 36+e^x }=\int\limits \frac{ 36+e^x-e^x }{ 36+e^x }dx=\int\limits 1dx-\int\limits \frac{ e^x }{ 36+e^x } dx=?\]